You've worked out a line of credit arrangement that allows you to borrow up to $35 million at any time. The interest rate is .52 percent per month. In addition, 2.5 percent of the amount that you borrow must be deposited in a non-interest-bearing account. Assume your bank uses compound interest on its line of credit loans. What is the effective annual interest rate on this lending arrangement?
Answer #1

Given,

Interest rate (r) = 0.52% or 0.0052

Non-interest bearing deposit rate = 2.5% or 0.025

Solution :-

EAR without compensating balance - EAR = (1 +8 12 months - I = (1 + 0.0057)K = (1,0052) - = 606422 -1 = 0.06472 og 6.427% EAR